show that [A2,B]=A[A,B]+[A,B]A


1
Expert's answer
2021-10-03T13:23:49-0400

We use the definition for the operator on a function f:


"[A^2,B]f=A[A,B]f+[A,B]Af
\\ [A^2,B]f=A^2Bf-BA^2f
\\ [A^2,B]f=A(AB)f-(BA)Af"


Following that, we add one term that is equal to zero to then rearrange and find the equivalence with the definitions for [A,B]:


"[A^2,B]f=A(AB)f-(BA)Af-A(BA)f+(AB)Af
\\ =A(AB)f-A(BA)f-(BA)Af+(AB)Af
\\ =A[AB-BA]f+[AB-BA]Af
\\ [A^2,B]f =A[A,B]f+[A,B]Af"


In conclusion, we were able to show that [A2,B]=A[A,B]+[A,B]A.



Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.